A company estimates that 0.2% of their products will fail after the original warranty period but within 2 years of the purchase, with a replacement cost of $500.


If they offer a 2 year extended warranty for $31, what is the company's expected value of each warranty sold?


1
Expert's answer
2021-02-24T07:48:20-0500

The outcomes are $"31" and -($"500" -$"31" )=-$"469"


The proabability of outcomes are "1-0.002=0.998 \text{ and }0.002" respectively.


The company's expected value of each warranty sold is-

                  

             "=31\times 0.998+(-469)\times 0.002=30"


Hence Company's expected value is $"30."